# Benchmarking Large Language Models for Math Reasoning Tasks

### v1

@@ Abstract
Introduction
Increasing computational resources and the free accessibility to large amounts of textual data have boosted the developments of Large Language Models (LLMs) (Brown et al. 2020; OpenAI 2023a; Touvron et al. 2023; Gemini-Team et al. 2023) in recent years. Central to this progress is the transformer architecture (Vaswani et al. 2017) based on the self-attention mechanism, which superseded recurrent neural networks for natural language processing and led to a trend of extremely large and powerful models, mastering many text-based tasks, such as summarizing (Zhang et al. 2024; Goyal, Li, and Durrett 2022), translating (Jiao et al. 2023) or question answering (Chowdhery et al. 2023).
While these models perform remarkably well for many language tasks, solving mathematical word problems still presents a challenge for LLMs (Hendrycks et al. 2021; Imani, Du, and Shrivastava 2023; Kaddour et al. 2023). Such problems require not only a deep understanding of the task description but also sophisticated logical and mathematical reasoning to determine the solution path. Additionally, the computational demands of finding the result pose further challenges. Since foundational models are not specifically optimized for mathematical calculations, they often struggle even with elementary arithmetic (Cobbe et al. 2021). Figure 1 lists various answers from different models.
The ability to perform mathematical reasoning is crucial not only in educational settings, where LLMs could potentially serve as a technological foundations for personalized tutoring (Liang et al. 2023; Kasneci et al. 2023), but also in higher education, where it could assist with complex problem-solving and discovering new proofs (Wang et al. 2024). Moreover, in professional fields such as finance and economics, where mathematical problems are prevalent, these capabilities could have significant applications (Korinek 2023).Therefore, enhancing LLMs’ mathematical reasoning skills would benefit various applications, extending from classroom support to advanced research and industry.
To address this challenge, several algorithms and methods have been proposed to improve the mathematical reasoning capabilities of LLMs. Understanding the strengths and limitations of each algorithm is essential for making informed decisions and effectively applying them in different contexts. While previous research has often evaluated these methods on rather limited datasets (Qiao et al. 2023) or using similar methods (Luo et al. 2023), there remains a need for more comprehensive analysis. Addressing this research gap, our study presents an extensive benchmarking of seven advanced methods across five widely-used datasets. Additionally, whereas prior work has mainly focused on accuracy as the main performance metric, we also consider time and cost factors, for deploying these algorithms effectively.
@@ Introduction
To tackle these challenges, there are different ideas for improving the mathematical reasoning capabilities of LLMs. Keeping the resource consumption low, the prompt for the LLM can be adapted by providing various samples and initiating a step-by-step problem solving approach. The Chain-of-Thought (CoT) approach (Wei et al. 2022) as well as the Auto CoT (Zhang et al. 2023) exploit this idea. In a similar way, the Zero-Shot CoT (Kojima et al. 2022) operates, but without the few-shot samples. Setting the focus on the robustness of the solution and therefore requiring a higher computational effort, the Self-Consistency (Wang et al. 2023) and Complex CoT (Fu et al. 2023) approaches optimize the solving process. Both call the foundational model repeatedly and therefore increase the robustness. Lastly, external engines like Python can be exploited outsourcing the computational tasks partly from the LLM to an external resource. PAL (Gao et al. 2023) and PoT (Chen et al. 2023) are both using this approach, applying few-shot samples and pipelines adapted for this use case. All these approaches bring their trade-off between accuracy, robustness, and consumption of resources, and for a fair comparison, all dimensions need to be investigated thoroughly.
@@
Related Work
Multiple surveys examine the problem of mathematical reasoning with LLMs from theoretical viewpoints. Lu et al. (2023) investigate the different deep learning methods used for mathematical reasoning, Ahn et al. (2024) give a thorough overview specifically covering LLMs for this task, and Huang and Chang (2023) analyse the different kinds of reasoning problems and the corresponding opportunities for LLMs. All surveys emphasize the missing generalizability and robustness, the challenges with complex questions, and the problem of hallucinations and trustworthiness.
Setting the focus on the algebraic skills of LLMs, Yuan et al. (2023) find that GPT-3.5 and GPT-4 outperform the other foundation models. As key points, they mark the influence of tokenization, the importance of the model size, and the sensitivity of prompts, identifying different optimal prompts across the LLMs. However, they also emphasize that the algebraic ability of a model cannot be directly equated with the mathematical reasoning capabilities.
From an empirical viewpoint Qiao et al. (2023) show that Chain-of-Thought reasoning works decently on the GSM8K dataset for GPT-3 (Brown et al. 2020) and Codex (Chen et al. 2021), but poor for LaMDA (Thoppilan et al. 2022), and the results for PaLM (Chowdhery et al. 2023) lie in-between. They compare different model sizes, in general, finding that bigger models perform better. Luo et al. (2023) compare various foundation models on the GSM8K and the MATH dataset, finding that closed-source models perform significantly better than open-source models and GPT-4 being the overall best model for mathematical reasoning.
A detailed analysis of various LLMs for mathematical reasoning tasks in combination with visual inputs is reported by Lu et al. (2024). In line with the previous results, they show the best performance for GPT-4 Vision (OpenAI 2023b), which lies still clearly below human achievements.
@@
Datasets for Mathematical Reasoning
To assess the reasoning ability of the methods, we use several mathematical word problems. These tasks stem from five publicly available datasets that are commonly used in the relevant literature and represent varying levels of difficulty, from elementary school to university mathematics. The datasets also include various types of problems, including both open-ended questions and multiple-choice tasks. The specific details of the datasets are presented in Table 1.
@@
Foundation Models
Many of the mathematical reasoning methods are model-agnostic, i.e. they are not dependent on a particular base model. Although most studies report results using GPT-3.5 and GPT-4, the choice of additional models varies. To ensure a fair and comprehensive comparison, we evaluate these methods using state-of-the-art open-source and closed-source base models. In addition to GPT-3.5 111gpt-3.5-turbo-0125 and GPT-4o 222gpt-4o-2024-05-13 (OpenAI 2023a), we analyze LLaMA 3-8B and LLaMA 3-70B models (Touvron et al. 2023).
Methods for Mathematical Reasoning
@@
The standard Chain-of-Thought (CoT) approach (Wei et al. 2022) involves providing the model with multiple examples where the solution process is detailed step by step, guiding the model to generate a reasoning chain before arriving at a conclusion. The Zero-Shot CoT approach (Kojima et al. 2022) streamlines this by replacing explicit examples with the key phrase "Let’s think step by step.". This strategy involves two model calls: one to generate the reasoning path and another to extract the final answer. A variant of the classic CoT is the Auto CoT approach (Zhang et al. 2023), which replaces manually curated few-shot examples with automatically selected ones. This is achieved by clustering the training data and selecting the best representative from each cluster.
Due to the stochastic nature of foundation models, the same prompt can yield both correct and incorrect solutions. To address this variability, the Self-Consistency CoT approach (Wang et al. 2023) involves generating multiple responses from the model and using a majority vote to select the most likely correct solution. This method utilizes the same few-shot examples as the standard CoT. In contrast, the Complex CoT approach (Fu et al. 2023) operates on the assumption that higher complexity can enhance performance. It selects the most complex examples from the training data for few-shot input and generates multiple solutions, applying majority voting only to the most complex responses.
Recognizing the limitations of LLMs in algebraic tasks, the Program-of-Thought (PoT) (Chen et al. 2023) approach extends CoT to make use of an external Python engine. In the few-shot examples, it is guided to separate the reasoning process from the computation part by generating program code. A similar idea is exploited by PAL (Gao et al. 2023), which integrates program synthesis into the model’s workflow, focusing on generating executable code to solve tasks.
@@
Performance Metrics
To evaluate the results, we use the pass@k metric (Chen et al. 2021), which measures the probability that the correct answer is generated at least once when the model is prompted times. Chen et al. (2021) introduced an unbiased estimator for this metric by prompting the model times, which helps mitigate the impact of the inherent stochasticity in language models. Given as the number of correct responses, the metric is calculated as follows:
@@
Robust Performance
In this section, we examine the influence of various prompt strategies, foundation models, and datasets on the pass@3 metric, which represents robust performance. Table 2 lists the comparison results. Besides the pass@3 metric, we also report the accuracy reached by all models in Appendix.
Foundation Models.
@@
Discussion
In our benchmarking study, we provide empirical evidence that the choice of the foundation model significantly influences the performance of LLMs on mathematical reasoning tasks, overshadowing the impact of specific prompting strategies. For instance, models like GPT-4o and LLaMA 3-70B show performance enhancements of up to 20% over smaller models across various datasets, highlighting the crucial role of model architecture and size. These findings align with previous results, indicating that model size significantly enhances the performance, and GPT-4 proving to be the strongest model (Qiao et al. 2023; Luo et al. 2023).
@@ Discussion
Despite these advances, there is still room for improvement in the performance of mathematical reasoning tasks at school and university level. For example, while GPT-4o achieves over 80% accuracy in the AQuA dataset, human performance is typically 85% according to (Zhong et al. 2023). In the MATH dataset, the maximum accuracy of GPT-4o is 75%, while PhD students achieve between 40% and 90% (Hendrycks et al. 2021). These results suggest that further progress in model training and reasoning skills is needed to close this gap, especially for tasks requiring deeper cognitive abilities and complex problem-solving.
@@ Conclusion
We identify the limited number of foundation models used in our benchmark. Future work should consider other state-of-the-art open-source models such as Mistral-AI (Jiang et al. 2023) and closed-source models like Gemini (Gemini-Team et al. 2023) or Claude (Anthropic 2024). Nevertheless, our codebase has been released and is easy to use for integrating new foundation models. Moreover, we identified another set of models fine-tuned on mathematical reasoning problems, which we plan to include in our benchmarking for a more comprehensive evaluation beyond just prompt strategies.
Appendix A Appendix
Proficiency Analysis
To further analyze the proficiency of GPT models on the MATH dataset, we compare the results across different levels and question types in Figures 6 and 7. GPT-3.5 demonstrates relatively high performance in Algebra and Pre-algebra questions, but for the remaining categories, its pass@3 metric falls below 50% at Levels 4 and 5. In contrast, GPT-4 shows strong performance across all types and levels, with only Level 5 questions in Geometry, Intermediate Algebra, and Pre-calculus scoring below 50% pass@3.
This proficiency can be observed in the following example of a Level 5 question from the Algebra category of the MATH dataset. Both GPT models successfully solve the problem, though GPT-4o again demonstrates its talkative nature requiring many tokens to arrive at the final result.
Restricted number of tokens
Across all datasets and methods, the GPT-4 model struggles with some questions, not due to poor reasoning quality, but because it requires too many tokens for its argumentation. The following example from the GSM8K dataset demonstrates how the model’s detailed thinking leads to exceeding the token limit before arriving at an answer when using the PoT method. For comparison, we have also included the GPT-3.5 output.
Statistical analysis
To test our hypothesis that the differences between the foundation models are significant, while between the methods they are not, we conduct an ANOVA analysis on the results of the MATH dataset. The analysis for the foundation models yields a p-value of 1.9e-5, indicating a highly significant difference between them. In contrast, the analysis for the methods resulted in a p-value of 0.72, suggesting that the differences between methods are not statistically significant.
We also conduct a more detailed analysis of the differences between the foundation models. As illustrated in Figure 8, the results reveal that the difference between GPT-3.5 and LLaMA 3-70B is the only comparison that is not statistically significant. This suggests that while most foundation models show significant performance variations, GPT-3.5 and LLaMA 3-70B are more comparable in their results on the MATH dataset. This finding further emphasizes the importance of choosing the right foundational model, as it has a substantial impact on performance.
Accuracy results
Table 3 lists all accuracy results, corresponding to the pass@1 metric. The distribution of these results is similar to the pass@3 metric. GPT-4o with Self-Consistency CoT consistently outperforms all other methods and foundational models.
GSM8K | SVAMP | Multi-Arith | AQuA | MATH | ||
CoT | GPT-3.5 | 0.60 0.35 | 0.73 0.34 | 0.95 0.15 | 0.54 0.36 | 0.41 0.40 |
GPT-4o | 0.93 0.21 | 0.94 0.22 | 1.00 0.00 | 0.75 0.34 | 0.71 0.40 | |
LLaMA 3-8B | 0.39 0.31 | 0.56 0.33 | 0.83 0.17 | 0.31 0.23 | 0.14 0.23 | |
LLaMA 3-70B | 0.67 0.32 | 0.75 0.28 | 0.92 0.10 | 0.34 0.30 | 0.23 0.29 | |
Auto CoT | GPT-3.5 | 0.47 0.33 | 0.83 0.31 | 0.77 0.27 | 0.58 0.37 | 0.36 0.38 |
GPT-4o | 0.88 0.24 | 0.94 0.21 | 0.99 0.04 | 0.65 0.33 | 0.70 0.39 | |
LLaMA 3-8B | 0.43 0.32 | 0.67 0.33 | 0.90 0.13 | 0.32 0.23 | 0.08 0.14 | |
LLaMA 3-70B | 0.78 0.30 | 0.84 0.26 | 0.99 0.05 | 0.53 0.33 | 0.31 0.35 | |
ZeroShot CoT | GPT-3.5 | 0.82 0.29 | 0.82 0.30 | 0.97 0.10 | 0.60 0.35 | 0.34 0.39 |
GPT-4o | 0.79 0.31 | 0.91 0.26 | 0.97 0.08 | 0.63 0.36 | 0.50 0.43 | |
LLaMA 3-8B | 0.31 0.25 | 0.46 0.28 | 0.56 0.23 | 0.10 0.11 | 0.06 0.14 | |
LLaMA 3-70B | 0.55 0.32 | 0.59 0.29 | 0.80 0.17 | 0.10 0.16 | 0.08 0.18 | |
Complex CoT | GPT-3.5 | 0.81 0.31 | 0.78 0.35 | 0.97 0.10 | 0.32 0.32 | 0.38 0.40 |
GPT-4o | 0.89 0.25 | 0.94 0.22 | 0.80 0.00 | 0.59 0.31 | 0.55 0.46 | |
LLaMA 3-8B | 0.36 0.34 | 0.64 0.36 | 0.93 0.13 | 0.38 0.26 | 0.14 0.24 | |
LLaMA 3-70B | 0.74 0.32 | 0.82 0.31 | 0.99 0.04 | 0.59 0.34 | 0.31 0.37 | |
SelfConsistency CoT | GPT-3.5 | 0.34 0.22 | 0.77 0.40 | 0.96 0.17 | 0.66 0.44 | 0.19 0.33 |
GPT-4o | 0.95 0.21 | 0.95 0.21 | 1.00 0.00 | 0.81 0.35 | 0.76 0.39 | |
LLaMA 3-8B | 0.31 0.21 | 0.72 0.37 | 1.00 0.01 | 0.45 0.33 | 0.18 0.31 | |
LLaMA 3-70B | 0.82 0.35 | 0.89 0.26 | 1.00 0.01 | 0.46 0.38 | 0.30 0.37 | |
PAL | GPT-3.5 | 0.68 0.40 | 0.76 0.40 | 0.92 0.27 | 0.30 0.40 | N/a |
GPT-4o | 0.88 0.25 | 0.89 0.23 | 0.90 0.15 | 0.36 0.37 | N/a | |
LLaMA 3-8B | 0.36 0.33 | 0.55 0.36 | 0.70 0.26 | 0.13 0.23 | N/a | |
LLaMA 3-70B | 0.65 0.36 | 0.72 0.35 | 0.88 0.23 | 0.22 0.33 | N/a | |
PoT | GPT-3.5 | 0.75 0.44 | 0.87 0.33 | 0.99 0.10 | 0.50 0.50 | N/a |
GPT-4o | 0.83 0.38 | 0.80 0.40 | 1.00 0.00 | 0.66 0.47 | N/a | |
LLaMA 3-8B | 0.54 0.50 | 0.58 0.49 | 0.42 0.49 | 0.31 0.46 | N/a | |
LLaMA 3-70B | 0.82 0.38 | 0.87 0.33 | 0.70 0.46 | 0.58 0.49 | N/a |
Visualized Efficiency Comparison
To balance performance and resource consumption, we present the results across all datasets visually in Figures 9, 10, 11 and 12. On the smallest dataset, Multiarith, which is at the grade school level, nearly all methods and models consistently achieve over 90% pass@3. Even the smaller LLaMA 3-8B model proves to be an efficient and cost-effective option, particularly when used with the AutoCoT method. For the SVAMP dataset, the results are similarly strong, with pass@3 rates consistently exceeding 70%. Although GPT-4o and LLaMA 3-70B slightly outperform the other models, the margin is small.
In contrast, the results for the more advanced, college-level AQuA dataset are more varied. Interestingly, GPT-3.5 with Zero-Shot CoT nearly matches the best performance of GPT-4o while being significantly more resource-efficient. For LLaMA 3-70B, AutoCoT and Complex CoT deliver the best results.
The MATH dataset reveals the largest disparity between open-source and closed-source models. While LLaMA 3 struggles to surpass the 50% mark, GPT-4o consistently exceeds 60%, and even GPT-3.5 achieves between 50% and 60% pass@3.
Few-shot examples
For reproducibility, we include all few-shot examples used in the benchmarking survey, which are also available in our GitHub repository. Merely Zero-Shot CoT did not use any specific examples but only the prompt Let’s think step by step. For CoT, we used the prompts from the original paper (Wei et al. 2022), and the same prompts were used for Self-Consistency, see Figures 13 and 14. For MATH dataset we used the first 8 samples from the according training dataset.
As suggested by the original paper, we extracted the samples for the Auto CoT method using a clustering approach (Zhang et al. 2023). All used samples can be found in Figures 22, 23, 24, 25, 26 and 27. The Complex CoT approach extracts the “most complex” samples from the training data, to guide the reasoning process (Fu et al. 2023). We report our extracted samples in Figures 28, 29, 30, 31, 32 and 33. Due to limited GPU space, we used only 6 shots for the MATH dataset in the Complex CoT approach.
For PAL, we use the prompts suggested by the original paper (Gao et al. 2023), as seen in Figure 15 and 16. PoT applies different prompts for all datasets (Chen et al. 2023), they can found in Figures 17, 18, 19, 20 and 21. For Multi-Arith dataset, the original implementation only offered a Zero-Shot Approach, which we applied.

### v2

@@ Abstract
Index Terms:
Large Language Models, Mathematical Reasoning, Generative AI, In-context Learning, Benchmarking StudyI Introduction
Increasing computational resources and the free accessibility to large amounts of textual data have boosted the developments of Large Language Models (LLMs) [1, 2, 3, 4] in recent years. Central to this progress is the transformer architecture [5] based on the self-attention mechanism, which superseded recurrent neural networks for natural language processing and led to a trend of extremely large and powerful models, mastering many text-based tasks, such as summarizing [6, 7], translating [8] or question answering [9].
While these models perform remarkably well for many language tasks, solving mathematical word problems still presents a challenge for LLMs [10, 11, 12]. Such problems require not only a deep understanding of the task description but also sophisticated logical and mathematical reasoning to determine the solution path. Additionally, the computational demands of finding the result pose further challenges. Since foundational models are not specifically optimized for mathematical calculations, they often struggle even with elementary arithmetic [13]. Figure 1 lists various answers from different models.
The ability to perform mathematical reasoning is crucial not only in educational settings, where LLMs could potentially serve as a technological foundations for personalized tutoring [14, 15], but also in higher education, where it could assist with complex problem-solving and discovering new proofs [16]. Moreover, in professional fields such as finance and economics, where mathematical problems are prevalent, these capabilities could have significant applications [17].Therefore, enhancing LLMs’ mathematical reasoning skills would benefit various applications, extending from classroom support to advanced research and industry.
To address this challenge, several algorithms and methods have been proposed to improve the mathematical reasoning capabilities of LLMs. Understanding the strengths and limitations of each algorithm is essential for making informed decisions and effectively applying them in different contexts. While previous research has often evaluated these methods on rather limited datasets [18] or using similar methods [19], there remains a need for more comprehensive analysis. Addressing this research gap, our study presents an extensive benchmarking of seven advanced methods across five widely-used datasets. Additionally, whereas prior work has mainly focused on accuracy as the main performance metric, we also consider time and cost factors, for deploying these algorithms effectively.
@@
To tackle these challenges, there are different ideas for improving the mathematical reasoning capabilities of LLMs. Keeping the resource consumption low, the prompt for the LLM can be adapted by providing various samples and initiating a step-by-step problem solving approach. The Chain-of-Thought (CoT) approach [20] as well as the Auto CoT [21] exploit this idea. In a similar way, the Zero-Shot CoT [22] operates, but without the few-shot samples. Setting the focus on the robustness of the solution and therefore requiring a higher computational effort, the Self-Consistency [23] and Complex CoT [24] approaches optimize the solving process. Both call the foundational model repeatedly and therefore increase the robustness. Lastly, external engines like Python can be exploited outsourcing the computational tasks partly from the LLM to an external resource. PAL [25] and PoT [26] are both using this approach, applying few-shot samples and pipelines adapted for this use case. All these approaches bring their trade-off between accuracy, robustness, and consumption of resources, and for a fair comparison, all dimensions need to be investigated thoroughly.
@@
II Related Work
Multiple surveys examine the problem of mathematical reasoning with LLMs from theoretical viewpoints. [27] investigate the different deep learning methods used for mathematical reasoning, [28] give a thorough overview specifically covering LLMs for this task, and [29] analyse the different kinds of reasoning problems and the corresponding opportunities for LLMs. All surveys emphasize the missing generalizability and robustness, the challenges with complex questions, and the problem of hallucinations and trustworthiness.
Setting the focus on the algebraic skills of LLMs, [30] find that GPT-3.5 and GPT-4 outperform the other foundation models. As key points, they mark the influence of tokenization, the importance of the model size, and the sensitivity of prompts, identifying different optimal prompts across the LLMs. However, they also emphasize that the algebraic ability of a model cannot be directly equated with the mathematical reasoning capabilities.
From an empirical viewpoint [18] show that Chain-of-Thought reasoning works decently on the GSM8K dataset for GPT-3 [1] and Codex [31], but poor for LaMDA [32], and the results for PaLM [9] lie in-between. They compare different model sizes, in general, finding that bigger models perform better. [19] compare various foundation models on the GSM8K and the MATH dataset, finding that closed-source models perform significantly better than open-source models and GPT-4 being the overall best model for mathematical reasoning.
A detailed analysis of various LLMs for mathematical reasoning tasks in combination with visual inputs is reported by [33]. In line with the previous results, they show the best performance for GPT-4 Vision [34], which lies still clearly below human achievements.
@@ III Benchmarking Details
III-A Datasets for Mathematical Reasoning
To assess the reasoning ability of the methods, we use several mathematical word problems. These tasks stem from five publicly available datasets that are commonly used in the relevant literature and represent varying levels of difficulty, from elementary school to university mathematics. The datasets also include various types of problems, including both open-ended questions and multiple-choice tasks. The specific details of the datasets are presented in Table I.
@@ III-A Datasets for Mathematical Reasoning
III-B Foundation Models
Many of the mathematical reasoning methods are model-agnostic, i.e. they are not dependent on a particular base model. Although most studies report results using GPT-3.5 and GPT-4, the choice of additional models varies. To ensure a fair and comprehensive comparison, we evaluate these methods using state-of-the-art open-source and closed-source base models. In addition to GPT-3.5 111gpt-3.5-turbo-0125 and GPT-4o 222gpt-4o-2024-05-13 [2], we analyze LLaMA 3-8B and LLaMA 3-70B models [3].
III-C Methods for Mathematical Reasoning
@@
The standard Chain-of-Thought (CoT) approach [20] involves providing the model with multiple examples where the solution process is detailed step by step, guiding the model to generate a reasoning chain before arriving at a conclusion. The Zero-Shot CoT approach [22] streamlines this by replacing explicit examples with the key phrase "Let’s think step by step.". This strategy involves two model calls: one to generate the reasoning path and another to extract the final answer. A variant of the classic CoT is the Auto CoT approach [21], which replaces manually curated few-shot examples with automatically selected ones. This is achieved by clustering the training data and selecting the best representative from each cluster.
Due to the stochastic nature of foundation models, the same prompt can yield both correct and incorrect solutions. To address this variability, the Self-Consistency CoT approach [23] involves generating multiple responses from the model and using a majority vote to select the most likely correct solution. This method utilizes the same few-shot examples as the standard CoT. In contrast, the Complex CoT approach [24] operates on the assumption that higher complexity can enhance performance. It selects the most complex examples from the training data for few-shot input and generates multiple solutions, applying majority voting only to the most complex responses.
Recognizing the limitations of LLMs in algebraic tasks, the Program-of-Thought (PoT) [26] approach extends CoT to make use of an external Python engine. In the few-shot examples, it is guided to separate the reasoning process from the computation part by generating program code. A similar idea is exploited by PAL [25], which integrates program synthesis into the model’s workflow, focusing on generating executable code to solve tasks.
@@
III-D Performance Metrics
To evaluate the results, we use the pass@k metric [31], which measures the probability that the correct answer is generated at least once when the model is prompted times. [31] introduced an unbiased estimator for this metric by prompting the model times, which helps mitigate the impact of the inherent stochasticity in language models. Given as the number of correct responses, the metric is calculated as follows:
@@ IV-A Experimental setup
IV-B Robust Performance
In this section, we examine the influence of various prompt strategies, foundation models, and datasets on the pass@3 metric, which represents robust performance. Table II lists the comparison results. Besides the pass@3 metric, we also report the accuracy reached by all models in Appendix.
IV-B1 Foundation Models.
@@ IV-D Proficiency
V Discussion
In our benchmarking study, we provide empirical evidence that the choice of the foundation model significantly influences the performance of LLMs on mathematical reasoning tasks, overshadowing the impact of specific prompting strategies. For instance, models like GPT-4o and LLaMA 3-70B show performance enhancements of up to 20% over smaller models across various datasets, highlighting the crucial role of model architecture and size. These findings align with previous results, indicating that model size significantly enhances the performance, and GPT-4 proving to be the strongest model [18, 19].
@@ V Discussion
Despite these advances, there is still room for improvement in the performance of mathematical reasoning tasks at school and university level. For example, while GPT-4o achieves over 80% accuracy in the AQuA dataset, human performance is typically 85% according to [35]. In the MATH dataset, the maximum accuracy of GPT-4o is 75%, while PhD students achieve between 40% and 90% [10]. These results suggest that further progress in model training and reasoning skills is needed to close this gap, especially for tasks requiring deeper cognitive abilities and complex problem-solving.
@@ VI Conclusion
We identify the limited number of foundation models used in our benchmark. Future work should consider other state-of-the-art open-source models such as Mistral-AI [36] and closed-source models like Gemini [4] or Claude [37]. Nevertheless, our codebase has been released and is easy to use for integrating new foundation models. Moreover, we identified another set of models fine-tuned on mathematical reasoning problems, which we plan to include in our benchmarking for a more comprehensive evaluation beyond just prompt strategies.